OpenStudy (anonymous):

When x = 2 and y = 3, z = 42. Write the function that models each of the following relationships. 1. z varies inversely with x and directly with y. z=y/x z=28y/x z=(28y)x z=3y/2x

OpenStudy (anonymous):

formula would be z = ky/x substitute: x = 2 y = 3 z = 42 formula becomes: 42 = k*3/2 solve for k to get: k = 2*42/3 = 28

OpenStudy (anonymous):

so it is z=(28y)x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

I got more cause I am not that good at math lol

OpenStudy (anonymous):

im not either

OpenStudy (anonymous):

well u helped me on this one lol